OpenStudy (anonymous):

Solve 2x2 - 8x = -7.

OpenStudy (kidrah69):

\[ax^2+bx+c\]

OpenStudy (kidrah69):

Put it in that form first then find A B C values

OpenStudy (anonymous):

A. \[-2\pm \sqrt{2}\] B. \[-2\pm2\sqrt{2}\] C. \[\frac{ 2\pm \sqrt{2} }{ 2 }\] D. \[2\pm \frac{ \sqrt{2} }{ 2 }\]

OpenStudy (welshfella):

you can use the quadratic formula x = [-b +/- sqrt(b^2 - 4ac)] / 2a

OpenStudy (anonymous):

A=2, B=-8, and C=7 right? @kidrah69

OpenStudy (welshfella):

plug values into the above formula for your equation b = -8, c = 7 and a = 2

OpenStudy (welshfella):

yes

OpenStudy (anonymous):

I got \[x = \frac{ -8\pm \sqrt{120} }{ 4 }\]

OpenStudy (anonymous):

@welshfella

OpenStudy (anonymous):

but the choices are more simplified

OpenStudy (welshfella):

-(-8) +/- sqrt(64-4*2*7) ------------------- 4

OpenStudy (welshfella):

= 8 +/- sqrt8 ---------- 4

OpenStudy (welshfella):

note that sqrt8 = 2sqrt2

OpenStudy (welshfella):

so we get 8 +/- 2 sqrt2 ---------- 4

OpenStudy (welshfella):

can you simplify that some more?

OpenStudy (anonymous):

So it's B?

OpenStudy (welshfella):

no divide 8 by 4 and 2 sqrt2 by 4

OpenStudy (welshfella):